the most recent financial statements for Assouad, Inc., are shown here: Income Statement Balance Sheet Sales $ 7,900 Current assets $ 3,900 Current liabilities $ 2,100 Costs 5,500 Fixed assets 8,600 Long-term debt 3,700 Taxable income $ 2,400 Equity 6,700 Taxes (25%) 600 Total $ 12,500 Total $ 12,500 Net income $ 1,800 Assets, costs, and current liabilities are proportional to sales. Long-term debt and equity are not. The company maintains a constant 40 percent dividend payout ratio. As with every other firm in its industry, next year’s sales are projected to increase by exactly 15 percent. What is the external financing needed? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1
Sales(7900*1.15) 9085
Costs(5500*1.15) 6325
Taxable income $2760
Taxes(25%*$2760) $690
Net income $2070
Less:Dividends($2070*40%) $828
Addition to retained earnings $1242

Total assets would be=$12500*1.15=$14375

Total equity would be=$6700+Addition to retained earnings

=(6700+1242)=$7942

Total current liabilities would be=$2100*1.15=$2415

Total liabilities would be=Current liabilities+Long term debt

=(2415+3700)=$6115

Total liabilities and equity=Total liabilities+Equity

=(6115+7942)=$14057

Total assets=Total liabilities+Total equity

Hence external financing needed=$14375-$14057

=$318.
Answer #2


Income statement % to sales Particulars Sales Cost Taxable income Tax at 25% Net Income Current year (Amount/sales) Next year